OpenStudy (anonymous):

using l'hopitale's rule solve lim x->0 (e^x+3x)^1/x

terenzreignz (terenzreignz):

You know the basics of L'Hôpital's rule?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

okay, so basically, if you have a limit \[\lim_{x \rightarrow a}\frac{f(x)}{g(x)}\] and\[f(x)\rightarrow0 \ and \ g(x)\rightarrow0\]OR \[f(x)\rightarrow \infty \ and \ g(x) \rightarrow \infty\] Then \[\lim_{x \rightarrow a}\frac{f(x)}{g(x)}=\lim_{x \rightarrow a}\frac{f \ '(x)}{g \ '(x)}\]

terenzreignz (terenzreignz):

Can you still follow?

terenzreignz (terenzreignz):

Sorry about that, my bad... Your question is of the form \[\huge \lim_{x \rightarrow a}f(x)^{g(x)}\] Where \[f(x) \rightarrow 1 \ and \ g(x) \rightarrow \infty\] Catch me so far?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Were you taught how to do this directly?

OpenStudy (anonymous):

yeah it was a while ago though

terenzreignz (terenzreignz):

Well, what were you taught? Or you want me to go over it step by step?

OpenStudy (anonymous):

to plug in the 0 for the values of x and if it was 0/0 then to find the derivative and plug in 0 again

terenzreignz (terenzreignz):

Well, that's not straightforward here, as you have a function raised to an exponent which happens to be another function, and not a function divided by another function... Let's go over how to do this, shall we? First, we assign a value to the limit, say \[\huge \lim_{x \rightarrow a}f(x)^{g(x)} = L\]

terenzreignz (terenzreignz):

Then let's apply the natural logarithm on both sides: \[\huge \ln \left(\lim_{x \rightarrow a}f(x)^{g(x)}\right)=\ln \ L\]

terenzreignz (terenzreignz):

The natural logarithm is a continuous function, so it can go inside the limit, like so: \[\huge \lim_{x \rightarrow a} \ \ln \left(f(x)^{g(x)}\right)=\ln \ L\]

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

Now, one property of the natural logarithm, a very useful one at that, is : \[\large \ln (a^b) = b \ln (a)\] Now to make good use of it...

terenzreignz (terenzreignz):

\[\huge \lim_{x \rightarrow a} \ g(x)\ln \left(f(x)\right)=\ln \ L\] Catch me so far?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

Well, however cumbersome, this can be rewritten as: \[\huge \lim_{x \rightarrow a}\frac{\ln (f(x))}{\frac{1}{g(x)}}=\ln \ L\]

terenzreignz (terenzreignz):

I need to know that you understand this step.

OpenStudy (anonymous):

i do

terenzreignz (terenzreignz):

Well, then, anyway.., we know that \[\large f(x) \rightarrow 1 \ and \ g(x) \rightarrow \infty\] In that case, I hope it's clear to you that \[\large \ln(f(x)) \rightarrow 0 \ and \frac1{g(x)} \rightarrow 0\] ... right?

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

Well, now you can use L'Hôpital's Rule on \[\huge \lim_{x \rightarrow a}\frac{\ln (f(x))}{\frac{1}{g(x)}}\] Since both the numerator and denominator go to zero. So you can evaluate its limit. However, that's not quite the end of it yet...

terenzreignz (terenzreignz):

Suppose ^that particular limit goes to some K. Then \[\huge \lim_{x \rightarrow a}\frac{\ln (f(x))}{\frac{1}{g(x)}} = \ln \ L = K\]

terenzreignz (terenzreignz):

Then, clearly \[\large L = e^K\]

terenzreignz (terenzreignz):

And it was L we wanted to find in the first place.

terenzreignz (terenzreignz):

So our final form is : \[\huge \lim_{x \rightarrow a}f(x)^{g(x)}=e^{\lim_{x \rightarrow a}\frac{\ln (f(x))}{\frac{1}{g(x)}}}\]

terenzreignz (terenzreignz):

That's not very neat, so let's clean it a bit: \[\huge \lim_{x \rightarrow a}f(x)^{g(x)} = e^K\]where \[\huge K= \lim_{x \rightarrow a}\frac{\ln (f(x))}{\frac{1}{g(x)}}\]

terenzreignz (terenzreignz):

And there you go; L'Hôpital's Rule for the case of the indeterminate form \[\huge 1^\infty\]

terenzreignz (terenzreignz):

Now, for your question, simply plug in \[f(x)=e^x + 3x\]\[g(x)=\frac{1}{x}\]

OpenStudy (anonymous):

okay thanks!